# Loss Gradient Gaussian Width based Generalization and Optimization Guarantees

### v1

@@ C.2 Gaussian Width bounds for Feed-Forward Networks (FFNs)
Note that the parameter vector $\mathbf{w}^{(l)}=\text{vec}(W^{(l)})$ and can be indexed with $j\in[m]$ and $j^{\prime}\in[d]$ when $l=1$ and $j^{\prime}\in[m]$ when $l\geq 2$. Then, we have
$\left[\frac{\partial\alpha^{(l)}}{\partial\mathbf{w}^{(l)}}\right]_{i,jj^{\prime}}$ $=\left[\frac{\partial\alpha^{(l)}}{\partial W^{(l)}}\right]_{i,jj^{\prime}}=\frac{1}{\sqrt{m}}\phi^{\prime}(\tilde{\alpha}^{(l)}_{i})\alpha^{(l-1)}_{j^{\prime}}\mathbbm{1}_{[i=j]}~{}.$ (99)
For $l\in\{2,\dots,L\}$, noting that $\frac{\partial\alpha^{(l)}}{\partial\mathbf{w}^{(l)}}\in\mathbb{R}^{m\times m^{2}}$ and $\left\lVert V\right\rVert_{F}=\left\lVert\operatorname{vec}(V)\right\rVert_{2}$ for any matrix $V$, we have
$\left\|\frac{\partial\alpha^{(l)}}{\partial\mathbf{w}^{(l)}}\right\|_{2}^{2}$ $=\sup_{\|V\|_{F}=1}\frac{1}{m}\sum_{i=1}^{m}\left(\phi^{\prime}(\tilde{\alpha}_{i}^{(l)})\sum^{m}_{j,j^{\prime}=1}\alpha^{(l-1)}_{j^{\prime}}\mathbbm{1}_{[i=j]}V_{jj^{\prime}}\right)^{2}$
$\leq\sup_{\|V\|_{F}=1}\frac{1}{m}\|V\alpha^{(l-1)}\|_{2}^{2}$
$\leq\frac{1}{m}\sup_{\|V\|_{F}=1}\|V\|_{2}^{2}\|\alpha^{(l-1)}\|_{2}^{2}$
$\overset{(a)}{\leq}\frac{1}{m}\|\alpha^{(l-1)}\|_{2}^{2}$
$\overset{(b)}{\leq}\frac{1}{m}\beta^{2(l-1)}~{},$
@@
$\left\|\frac{\partial\alpha^{(1)}}{\partial\mathbf{w}^{(1)}}\right\|_{2}^{2}\leq\frac{1}{m}\|\alpha^{(0)}\|_{2}^{2}=\frac{1}{m}\left\lVert\mathbf{x}\right\rVert_{2}^{2}=\frac{1}{m}~{},$
@@
We now restate and prove the main result for FFNs. See 5
@@
$f(\theta;\mathbf{x})=\mathbf{v}^{\top}\phi(\frac{1}{\sqrt{m}}W^{(L)}\phi(\cdots\phi(\frac{1}{\sqrt{m}}W^{(1)}\mathbf{x}))))~{},$ (103)
@@
$\alpha^{(l)}$ $=\phi\left(\frac{1}{\sqrt{m}}W^{(l)}\alpha^{(l-1)}(\mathbf{x})\right)$ (106)
@@
which is the output of the last layer or the so-called featurizer, i.e., $h^{(L)}(W,\mathbf{x})=\alpha^{(L)}(\mathbf{x})$. Then, the Gaussian width is simply $w(\hat{A}_{n}^{(L)})$.
For any hidden layer parameter $\mathbf{w}^{(l)}\in\mathbb{R}^{m^{2}}$ with $\mathbf{w}^{(l)}=\operatorname{vec}(W^{(l)})$, the gradient is in $\mathbb{R}^{m^{2}}$ and we have
@@
Define $\mathcal{Z}^{(l)}=\left\{Z^{(l)}\in\mathbb{R}^{m^{2}\times m}:\left\|Z^{(l)}\right\|_{2}\leq\frac{\beta^{l-1}}{\sqrt{m}}\right\}$, then according to Lemma 7, $\frac{\partial\alpha^{(l)}}{\partial\mathbf{w}^{(l)}}\in\mathcal{Z}^{(l)}$ with probability $1-\frac{2l}{m}$. Similarly, define =$\mathcal{B}=\left\{B\in\mathbb{R}^{m\times m}:\left\|B\right\|_{2}\leq\beta\right\}$, then according to Lemma 97, $\frac{\partial\alpha^{(l^{\prime})}}{\partial\alpha^{(l^{\prime}-1)}}\in\mathcal{B}$ with probability $1-\frac{2}{m}$. Besides, define $\mathcal{V}$ the set of $\mathbf{v}$, then according to (95), $\mathbf{v}\in\mathbb{R}^{m}$ and $\left\|\mathbf{v}\right\|_{2}\leq 1+\rho_{1}$, so $w\left(\mathcal{V}\right)\leq c\sup_{\mathbf{v}\in\mathcal{V}}\left\|\mathbf{v}\right\|_{2}\sqrt{m}\leq c(1+\rho_{1})\sqrt{m}$.
@@
$w(\mathcal{H}^{(l)})$ $\overset{(a)}{\leq}C\sup_{Z^{(l)}\in\mathcal{Z}^{(l)}}\|Z^{(l)}\|_{2}\prod_{l^{\prime}=l+1}^{L}\sup_{B\in\mathcal{B}}\|B\|_{2}~{}w(\mathcal{V})$
$\overset{(b)}{\leq}C\frac{\beta^{l-1}}{\sqrt{m}}\beta^{L-l}c\left(1+\rho_{1}\right)\sqrt{m}$
$=C_{0}(1+\rho_{1})\beta^{L-1}~{}.$
where (a) follows from Lemma 9 and (b) from the definition of $\mathcal{Z}^{(l)}$ and $\mathcal{B}$.
Define ^$\hat{\mathcal{H}}_{n}^{(l)}=\left\{\frac{1}{n}\sum_{z\in z^{(n)}}\frac{\partial f}{\partial\mathbf{w}^{(l)}}(z)\mid W\in B_{\rho}^{\operatorname{Spec}}(\theta_{0})\right\}$, from a similar proof as (75) we have
$w(\hat{\mathcal{H}}_{n}^{(l)})\leq w(\mathcal{H}^{(l)})\leq C_{0}(1+\rho_{1})\beta^{L-1}$
@@
According to Lemma 8, the Gaussian width of the average predictor gradient set is bounded by $\sum_{l=1}^{L}w(\hat{\mathcal{H}}_{n}^{(l)})+w(\hat{A}_{n}^{(L)})\leq C_{0}L(1+\rho_{1})\beta^{L-1}+w(\hat{A}_{n}^{(L)})$, then there exists constant $c_{1},c_{2}$ such that the Gaussian width of the loss gradient set is bounded by $c_{1}w(\hat{A}_{n}^{(L)})+c_{2}L(1+\rho_{1})\beta^{L-1}$. Therefore,
$w(\hat{\Xi}_{n}^{\text{ffn}})\leq c_{1}w(\hat{A}_{n}^{(L)})+c_{2}L(1+\rho_{1})\beta^{L-1}$
Since the spectral norm bound for each layer except the last layer holds with probability $(1-2/m)$ according to Lemma 4, and we use it $L$ times in the proof, the Gaussian width bound holds with probability $(1-2L/m)$. ∎
C.3 Gaussian Width bounds for Residual Networks (ResNets)
@@
$\begin{split}\alpha^{(0)}(\mathbf{x})&=\mathbf{x}~{},\\ \alpha^{(l)}(\mathbf{x})&=\alpha^{(l-1)}+\phi\left(\frac{1}{L\sqrt{m}}W^{(l)}\alpha^{(l-1)}(\mathbf{x})\right)~{},~{}~{}~{}l=1,\ldots,L~{},\\ f(\theta;\mathbf{x})=\alpha^{(L+1)}(\mathbf{x})&=\mathbf{v}^{\top}\alpha^{(L)}(\mathbf{x})~{},\end{split}$ (110)
@@
Consider any $l\in[L]$ and let $\beta:=\sigma_{1}+\frac{\rho}{\sqrt{m}}$. Under Assumptions 4 and 5, for $\theta\in B_{\rho,\rho_{1}}^{\operatorname{Spec}}(\theta_{0})$, with probability at least $\left(1-\frac{2l}{m}\right)$, we have
$\|\alpha^{(l)}\|_{2}\leq\left(1+\frac{\beta}{L}\right)^{l}~{}.$ (112)
@@
$\left\|\phi\left(\frac{1}{L\sqrt{m}}W^{(1)}\alpha^{(0)}\right)\right\|_{2}-\|\phi(\mathbf{0})\|_{2}\leq\left\|\phi\left(\frac{1}{L\sqrt{m}}W^{(1)}\alpha^{(0)}\right)-\phi(\mathbf{0})\right\|_{2}\leq\left\|\frac{1}{L\sqrt{m}}W^{(1)}\alpha^{(0)}\right\|_{2}~{},$
@@
$\|\alpha^{(1)}\|_{2}$ $\leq\|\alpha^{(0)}\|_{2}+\left\|\phi\left(\frac{1}{L\sqrt{m}}W^{(1)}\alpha^{(0)}\right)\right\|_{2}$
$\leq 1+\left\|\frac{1}{L\sqrt{m}}W^{(1)}\alpha^{(0)}\right\|_{2}$
$\leq 1+\frac{1}{L\sqrt{m}}\|W^{(1)}\|_{2}\|\alpha^{(0)}\|_{2}$
$\overset{(a)}{\leq}1+\frac{\beta}{L}~{},$
where (a) follows from Lemma 4 which holds with probability at least $1-\frac{2}{m}$.
@@
$\|\alpha^{(l-1)}\|_{2}\leq\left(1+\frac{\beta}{L}\right)^{l-1}~{},$
which holds with the probability at least $1-\frac{2(l-1)}{m}$. Since $\phi$ is 1-Lipschitz, for layer $l$, we have
$\left\|\phi\left(\frac{1}{L\sqrt{m}}W^{(l)}\alpha^{(l-1)}\right)\right\|_{2}-\|\phi(\mathbf{0})\|_{2}$ $\leq\left\|\phi\left(\frac{1}{L\sqrt{m}}W^{(l)}\alpha^{(l-1)}\right)-\phi(\mathbf{0})\right\|_{2}\leq\left\|\frac{1}{L\sqrt{m}}W^{(l)}\alpha^{(l-1)}\right\|_{2}~{},$
@@
$\|\alpha^{(l)}\|_{2}$ $\leq\|\alpha^{(l-1)}\|_{2}+\left\|\phi\left(\frac{1}{L\sqrt{m}}W^{(l)}\alpha^{(l-1)}\right)\right\|_{2}$
$\leq\|\alpha^{(l-1)}\|_{2}+\left\|\frac{1}{\sqrt{m}}W^{(l)}\alpha^{(l-1)}\right\|_{2}$
$\leq\|\alpha^{(l-1)}\|_{2}+\frac{1}{L\sqrt{m}}\|W^{(l)}\|_{2}\|\alpha^{(l-1)}\|_{2}$
$\overset{(a)}{\leq}\|\alpha^{(l-1)}\|_{2}+\frac{\beta}{L}\|\alpha^{(l-1)}\|_{2}$
$=\left(1+\frac{\beta}{L}\right)\|\alpha^{((l-1)}\|_{2}$
$\overset{(b)}{\leq}\left(1+\frac{\beta}{L}\right)^{l}~{},$
where (a) follows from Lemma 4 and (b) folows from the inductive step. Since we have used Lemma 4 $l$ times, after a union bound, our result would hold with probability at least $1-\frac{2l}{m}$. That completes the proof. ∎
@@
$\alpha^{(l)}=\alpha^{(l-1)}+\phi\left(\tilde{\alpha}^{(l)}\right)~{},~{}~{}~{}\tilde{\alpha}^{(l)}:=\frac{1}{L\sqrt{m_{l-1}}}W^{(l)}\alpha^{(l-1)}~{}.$ (113)
@@
Consider any $l\in\{2,\dots,L\}$ and let $\beta:=\sigma_{1}+\frac{\rho}{\sqrt{m}}$. Under Assumptions 4 and 5, for $\theta\in B_{\rho,\rho_{1}}^{\operatorname{Spec}}(\theta_{0})$, with probability at least $\left(1-\frac{2}{m}\right)$,
$\left\|\frac{\partial\alpha^{(l)}}{\partial\alpha^{(l-1)}}\right\|_{2}\leq 1+\frac{\beta}{L}~{}.$ (114)
@@
$\left[\frac{\partial\alpha^{(l)}}{\partial\alpha^{(l-1)}}\right]_{i,j}=\mathbbm{1}_{[i=j]}+\frac{1}{L\sqrt{m}}\phi^{\prime}(\tilde{\alpha}^{(l)}_{i})W_{ij}^{(l)}~{}.$ (115)
@@
$\left\|C^{(l)}\right\|_{2}^{2}$ $=\sup_{\|\mathbf{u}\|_{2}=1}\frac{1}{L^{2}m}\sum_{i=1}^{m}\left(\phi^{\prime}(\tilde{\alpha}^{(l)}_{i})\sum^{m}_{j=1}W_{ij}^{(l)}u_{j}\right)^{2}$
$\overset{(a)}{\leq}\sup_{\|\mathbf{u}\|_{2}=1}\frac{1}{L^{2}m}\|W^{(l)}\mathbf{u}\|_{2}^{2}$
$=\frac{1}{L^{2}m}\|W^{(l)}\|_{2}^{2}$
$\overset{(b)}{\leq}\left(\frac{\beta}{L}\right)^{2}~{},$
@@
Consider any $l\in[L]$. Under Assumptions 4 and 5, for $\theta\in B_{\rho,\rho_{1}}^{\operatorname{Spec}}(\theta_{0})$, with probability at least $\left(1-\frac{2l}{m}\right)$,
$\begin{split}\left\|\frac{\partial\alpha^{(l)}}{\partial\mathbf{w}^{(l)}}\right\|_{2}&\leq\frac{1}{L\sqrt{m}}\left(1+\frac{\beta}{L}\right)^{l-1}~{}.\end{split}$ (117)

### v2

@@ C.1 Gaussian Width bounds for Feed-Forward Networks (FFNs)
Note that the parameter vector $\mathbf{w}^{(l)}=\text{vec}(W^{(l)})$ and can be indexed with $j\in[m_{1}]$ and $j^{\prime}\in[d]$ when $l=1$, $j\in[m_{l}]$ and $j^{\prime}\in[m_{l-1}]$ when $l\geq 2$. Then, we have
$\left[\frac{\partial\alpha^{(l)}}{\partial\mathbf{w}^{(l)}}\right]_{i,jj^{\prime}}$ $=\left[\frac{\partial\alpha^{(l)}}{\partial W^{(l)}}\right]_{i,jj^{\prime}}=\frac{1}{\sqrt{m_{l}}}\phi^{\prime}(\tilde{\alpha}^{(l)}_{i})\alpha^{(l-1)}_{j^{\prime}}\mathbbm{1}_{[i=j]}~{}.$ (86)
For $l\in\{2,\dots,L\}$, noting that $\frac{\partial\alpha^{(l)}}{\partial\mathbf{w}^{(l)}}\in\mathbb{R}^{m_{l}\times m_{l}m_{l-1}}$ and $\left\lVert V\right\rVert_{F}=\left\lVert\operatorname{vec}(V)\right\rVert_{2}$ for any matrix $V$, we have
$\left\|\frac{\partial\alpha^{(l)}}{\partial\mathbf{w}^{(l)}}\right\|_{2}^{2}$ $=\sup_{\|V\|_{F}=1}\frac{1}{m_{l}}\sum_{i=1}^{m_{l}}\left(\phi^{\prime}(\tilde{\alpha}_{i}^{(l)})\sum_{j=1}^{m_{l}}\sum_{j^{\prime}=1}^{m_{l-1}}\alpha^{(l-1)}_{j^{\prime}}\mathbbm{1}_{[i=j]}V_{jj^{\prime}}\right)^{2}$
$\leq\sup_{\|V\|_{F}=1}\frac{1}{m_{l}}\|V\alpha^{(l-1)}\|_{2}^{2}$
$\leq\frac{1}{m_{l}}\sup_{\|V\|_{F}=1}\|V\|_{2}^{2}\|\alpha^{(l-1)}\|_{2}^{2}$
$\overset{(a)}{\leq}\frac{1}{m_{l}}\|\alpha^{(l-1)}\|_{2}^{2}$
$\overset{(b)}{\leq}\frac{1}{m_{l}}\prod_{k=1}^{l-1}\beta_{k}^{2}~{},$
@@
$\left\|\frac{\partial\alpha^{(1)}}{\partial\mathbf{w}^{(1)}}\right\|_{2}^{2}\leq\frac{1}{m_{1}}\|\alpha^{(0)}\|_{2}^{2}=\frac{1}{m_{1}}\left\lVert\mathbf{x}\right\rVert_{2}^{2}=\frac{1}{m_{1}}~{},$
@@
We now restate and prove the main result for FFNs. See 4
@@
$f(\theta;\mathbf{x})=\mathbf{v}^{\top}\phi(\frac{1}{\sqrt{m_{L}}}W^{(L)}\phi(\cdots\phi(\frac{1}{\sqrt{m_{1}}}W^{(1)}\mathbf{x}))))~{},$ (90)
@@
$\alpha^{(l)}$ $=\phi\left(\frac{1}{\sqrt{m_{l}}}W^{(l)}\alpha^{(l-1)}(\mathbf{x})\right)$ (93)
@@
which is the output of the last layer or the so-called featurizer, i.e., $h^{(L)}(W,\mathbf{x})=\alpha^{(L)}(\mathbf{x})$. Then, the Gaussian width is simply $w(A^{(L)})$.
For any hidden layer parameter $\mathbf{w}^{(l)}\in\mathbb{R}^{m_{l}m_{l-1}}$ with $\mathbf{w}^{(l)}=\operatorname{vec}(W^{(l)})$, the gradient is in $\mathbb{R}^{m_{l}m_{l-1}}$ and we have
@@
Define $\mathcal{Z}^{(l)}=\left\{Z^{(l)}\in\mathbb{R}^{m_{l}\times m_{l}m_{l-1}}:\left\|Z^{(l)}\right\|_{2}\leq\frac{\prod_{k=1}^{l-1}\beta_{k}}{\sqrt{m_{l}}}\right\}$, then according to Lemma 9, $\frac{\partial\alpha^{(l)}}{\partial\mathbf{w}^{(l)}}\in\mathcal{Z}^{(l)}$ with probability $1-\sum_{k=1}^{l-1}\frac{2}{m_{k}}$. Similarly, define $\mathcal{B}^{(l)}=\left\{B^{(l)}\in\mathbb{R}^{m_{l}\times m_{l-1}}:\left\|B^{(l)}\right\|_{2}\leq\beta_{l}\right\}$, then according to Lemma 84, $\frac{\partial\alpha^{(l)}}{\partial\alpha^{(l-1)}}\in\mathcal{B}^{(l)}$ with probability $1-\frac{2}{m_{l}}$. Besides, define $\mathcal{V}$ the set of $\mathbf{v}$, then according to (82), $\mathbf{v}\in\mathbb{R}^{m_{L}}$ and $\left\|\mathbf{v}\right\|_{2}\leq 1+\rho_{1}$, so $w\left(\mathcal{V}\right)\leq c\sup_{\mathbf{v}\in\mathcal{V}}\left\|\mathbf{v}\right\|_{2}\sqrt{m_{L}}\leq c(1+\rho_{1})\sqrt{m_{L}}$.
@@
$w(\mathcal{H}^{(l)})$ $\overset{(a)}{\leq}C\sup_{Z^{(l)}\in\mathcal{Z}^{(l)}}\|Z^{(l)}\|_{2}\prod_{l^{\prime}=l+1}^{L}\sup_{B^{(l^{\prime})}\in\mathcal{B}^{l^{\prime}}}\|B^{(l^{\prime})}\|_{2}~{}w(\mathcal{V})$
$\overset{(b)}{\leq}C\frac{\prod_{k=1}^{l-1}\beta_{k}}{\sqrt{m_{l}}}\prod_{k=l+1}^{L}\beta_{k}\cdot c\left(1+\rho_{1}\right)\sqrt{m_{L}}$
=$C_{0}\sqrt{\frac{m_{L}}{m_{l}}}(1+\rho_{1})\prod_{k\not=l}\beta_{k}~{}.$
where (a) follows from Lemma 11 and (b) from the definition of $\mathcal{Z}^{(l)}$ and $\mathcal{B}^{(l)}$.
@@
According to Lemma 10, the Gaussian width of the individual predictor gradient set is bounded by
$\sum_{l=1}^{L}w(\mathcal{H}^{(l)})+w(A^{(L)})$ $\leq\sum_{l=1}^{L}C_{0}\sqrt{\frac{m_{L}}{m_{l}}}(1+\rho_{1})\prod_{k\not=l}\beta_{k}+w(A^{(L)})$
$=C_{0}(1+\rho_{1})\sqrt{m_{L}}\prod_{l=1}^{L}\beta_{l}\sum_{l=1}^{L}\frac{1}{\beta_{l}\sqrt{m_{l}}}+w(A^{(L)})$
then there exists constant $c_{1},c_{2}$ such that
$w(\Xi^{\text{ffn}})\leq c_{1}w(A^{(L)})+c_{2}(1+\rho_{1})\sqrt{m_{L}}\prod_{l=1}^{L}\beta_{l}\sum_{l=1}^{L}\frac{1}{\beta_{l}\sqrt{m_{l}}}$
Since the spectral norm bound for each layer except the last layer holds with probability $(1-\frac{2}{m_{l}})$ according to Lemma 6, and we use it $L$ times in the proof, the Gaussian width bound holds with probability $(1-\sum_{l=1}^{L}\frac{2}{m_{l}})$. ∎
C.2 Gaussian Width bounds for Residual Networks (ResNets)
@@
$\begin{split}\alpha^{(0)}(\mathbf{x})&=\mathbf{x}~{},\\ \alpha^{(l)}(\mathbf{x})&=\alpha^{(l-1)}(\mathbf{x})+\phi\left(\frac{1}{L\sqrt{m_{l}}}W^{(l)}\alpha^{(l-1)}(\mathbf{x})\right)~{},~{}~{}~{}l=1,\ldots,L~{},\\ f(\theta;\mathbf{x})=\alpha^{(L+1)}(\mathbf{x})&=\mathbf{v}^{\top}\alpha^{(L)}(\mathbf{x})~{},\end{split}$ (97)
@@
Consider any $l\in[L]$ and let $\beta:=\sigma_{1}+\frac{\rho}{\sqrt{m}}$. Under Assumptions 3 and 4, for $\theta\in B_{\rho,\rho_{1}}^{\operatorname{Spec}}(\theta_{0})$, with probability at least $\left(1-\sum_{k=1}^{l}\frac{2}{m_{k}}\right)$, we have
$\|\alpha^{(l)}\|_{2}\leq\prod_{k=1}^{l}\left(1+\frac{\beta_{k}}{L}\right)~{}.$ (99)
@@
$\left\|\phi\left(\frac{1}{L\sqrt{m_{1}}}W^{(1)}\alpha^{(0)}\right)\right\|_{2}-\|\phi(\mathbf{0})\|_{2}\leq\left\|\phi\left(\frac{1}{L\sqrt{m_{1}}}W^{(1)}\alpha^{(0)}\right)-\phi(\mathbf{0})\right\|_{2}\leq\left\|\frac{1}{L\sqrt{m_{1}}}W^{(1)}\alpha^{(0)}\right\|_{2}~{},$
@@
$\|\alpha^{(1)}\|_{2}$ $\leq\|\alpha^{(0)}\|_{2}+\left\|\phi\left(\frac{1}{L\sqrt{m_{1}}}W^{(1)}\alpha^{(0)}\right)\right\|_{2}$
$\leq 1+\left\|\frac{1}{L\sqrt{m_{1}}}W^{(1)}\alpha^{(0)}\right\|_{2}$
$\leq 1+\frac{1}{L\sqrt{m_{1}}}\|W^{(1)}\|_{2}\|\alpha^{(0)}\|_{2}$
$\overset{(a)}{\leq}1+\frac{\beta_{1}}{L}~{},$
where (a) follows from Lemma 6 which holds with probability at least $1-\frac{2}{m_{1}}$.
@@
$\|\alpha^{(l-1)}\|_{2}\leq\prod_{k=1}^{l-1}\left(1+\frac{\beta_{k}}{L}\right)~{},$
which holds with the probability at least $1-\sum_{k=1}^{l-1}\frac{2}{m_{k}}$. Since $\phi$ is 1-Lipschitz, for layer $l$, we have
$\left\|\phi\left(\frac{1}{L\sqrt{m_{l}}}W^{(l)}\alpha^{(l-1)}\right)\right\|_{2}-\|\phi(\mathbf{0})\|_{2}$ $\leq\left\|\phi\left(\frac{1}{L\sqrt{m_{l}}}W^{(l)}\alpha^{(l-1)}\right)-\phi(\mathbf{0})\right\|_{2}\leq\left\|\frac{1}{L\sqrt{m_{l}}}W^{(l)}\alpha^{(l-1)}\right\|_{2}~{},$
@@
$\|\alpha^{(l)}\|_{2}$ $\leq\|\alpha^{(l-1)}\|_{2}+\left\|\phi\left(\frac{1}{L\sqrt{m_{l}}}W^{(l)}\alpha^{(l-1)}\right)\right\|_{2}$
$\leq\|\alpha^{(l-1)}\|_{2}+\left\|\frac{1}{L\sqrt{m_{l}}}W^{(l)}\alpha^{(l-1)}\right\|_{2}$
$\leq\|\alpha^{(l-1)}\|_{2}+\frac{1}{L\sqrt{m_{l}}}\|W^{(l)}\|_{2}\|\alpha^{(l-1)}\|_{2}$
$\overset{(a)}{\leq}\|\alpha^{(l-1)}\|_{2}+\frac{\beta_{l}}{L}\|\alpha^{(l-1)}\|_{2}$
$=\left(1+\frac{\beta_{l}}{L}\right)\|\alpha^{((l-1)}\|_{2}$
$\overset{(b)}{\leq}\prod_{k=1}^{l}\left(1+\frac{\beta_{k}}{L}\right)~{},$
where (a) follows from Lemma 6 and (b) follows from the inductive step. Since we have used Lemma 6 $l$ times, after a union bound, our result would hold with probability at least $1-\sum_{k=1}^{l}\frac{2}{m_{k}}$. That completes the proof. ∎
@@
$\alpha^{(l)}=\alpha^{(l-1)}+\phi\left(\tilde{\alpha}^{(l)}\right)~{},~{}~{}~{}\tilde{\alpha}^{(l)}:=\frac{1}{L\sqrt{m_{l}}}W^{(l)}\alpha^{(l-1)}~{}.$ (100)
@@
Consider any $l\in\{2,\dots,L\}$ and let $\beta_{l}:=\sigma_{1}+\frac{\rho}{\sqrt{m_{l}}}$. Under Assumptions 3 and 4, for $\theta\in B_{\rho,\rho_{1}}^{\operatorname{Spec}}(\theta_{0})$, with probability at least $\left(1-\frac{2}{m_{l}}\right)$,
$\left\|\frac{\partial\alpha^{(l)}}{\partial\alpha^{(l-1)}}\right\|_{2}\leq 1+\frac{\beta_{l}}{L}~{}.$ (101)
@@
$\left[\frac{\partial\alpha^{(l)}}{\partial\alpha^{(l-1)}}\right]_{i,j}=\mathbbm{1}_{[i=j]}+\frac{1}{L\sqrt{m_{l}}}\phi^{\prime}(\tilde{\alpha}^{(l)}_{i})W_{ij}^{(l)}~{}.$ (102)
@@
$\left\|C^{(l)}\right\|_{2}^{2}$ $=\sup_{\|\mathbf{u}\|_{2}=1}\frac{1}{L^{2}m_{l}}\sum_{i=1}^{m_{l}}\left(\phi^{\prime}(\tilde{\alpha}^{(l)}_{i})\sum_{j=1}^{m_{l-1}}W_{ij}^{(l)}u_{j}\right)^{2}$
$\overset{(a)}{\leq}\sup_{\|\mathbf{u}\|_{2}=1}\frac{1}{L^{2}m_{l}}\|W^{(l)}\mathbf{u}\|_{2}^{2}$
$=\frac{1}{L^{2}m_{l}}\|W^{(l)}\|_{2}^{2}$
$\overset{(b)}{\leq}\left(\frac{\beta_{l}}{L}\right)^{2}~{},$
@@
Consider any $l\in[L]$. Under Assumptions 3 and 4, for $\theta\in B_{\rho,\rho_{1}}^{\operatorname{Spec}}(\theta_{0})$, with probability at least $\left(1-\sum_{k=1}^{l}\frac{2}{m_{k}}\right)$,
$\begin{split}\left\|\frac{\partial\alpha^{(l)}}{\partial\mathbf{w}^{(l)}}\right\|_{2}&\leq\frac{1}{L\sqrt{m_{l}}}\prod_{k=1}^{l-1}\left(1+\frac{\beta_{k}}{L}\right)~{}.\end{split}$ (104)